# R-SFLLM: Jamming Resilient Framework for Split Federated Learning with Large Language Models

### v2

@@ 13             Set the user scheduling to $\alpha_{qnk}=1$ for all $(n,k)\in\mathcal{I}_{q}$ and 0 otherwise
III-C Iterative Water-Filling Solution
We have further shown in [22] that the NP-hard problem in (36) can be iteratively solved using a water-filling approach, as described in Algorithm 1. The proposed method adapts the original water-filling for MAC and MIMO channels [35] to incorporate user scheduling. To this end, each user $q\in\mathcal{Q}$ determines its update on the matrix $\boldsymbol{X}_{qnk}$ and computes an optimal set of the wireless system design parameters $\alpha_{qnk}$, $p_{qnk}$, $\boldsymbol{w}_{qnk}$, as outlined in steps (6)-(11), where for each user the following optimization problem is solved:
@@ III-C Iterative Water-Filling Solution
Algorithm 1 effectively circumvents the need to deal with the NP-hardness of the problem as the overall sum rate is indirectly maximized by maximizing the sum rate $R_{q}$ of the strongest users individually. In each iteration, we use the power iteration method for computing eigenvalues and eigenvectors, and perform water-filling for power allocation. Both of these methods are known to exhibit rapid convergence. Consequently, Algorithm 1 inherits these convergence properties, which we have verified in extensive experiments, where our method on average takes less than five iterations to converge. With $n_{\text{iter}}$ representing the number of iterations required for convergence, the overall complexity of the algorithm is
@@
III-D Worst-Case Jamming Strategy
In Proposition 2, we have seen that jamming LLM word embeddings affects the training performance. However, the impact still remains to be investigated for worst-case conditions and in particular how this affects the global performance in SFL after aggregating such corrupted models. To this end, we need to derive the worst-case jamming strategy, which we use to benchmark R-SFLLM. In contrast to anti-jamming, we need to find the covariance matrix $\boldsymbol{C}_{\boldsymbol{u}_{nk}}$, which minimizes $R$ instead, and pose the following adversarial objective:
@@ III-D Worst-Case Jamming Strategy
This is a convex semidefinite program due to the objective function and the constraints being convex. Thus, a global optimum can be found in polynomial time using interior-point or first-order methods [36]. However, neither one of these approaches scales efficiently for high-dimensions with a complexity of $\mathcal{O}(N_{J}^{4}N^{2}K^{2})$, or higher [37]. Thus, we require a compute-efficient alternative. To this end, we proposed a two-step approximation procedure in [22] described in Algorithm 2, which consists of a prior user selection stage and a subsequent compute-efficient convex optimization. In [22], we have further shown that this worst-case jammer nullifies the sum rate, independent of the number of antennas and DoAs. Thus, we adopt this adversary as a benchmark in the subsequent experiments and investigate its impact on the global model performance in SFL with LLMs to answer the question whether worst-case jamming translates into worst-case SFL training performance.
@@
4for each user $q\in\mathcal{Q}$ do
@@
8Determine the strongest user $q^{*}=\arg\max_{q\in\mathcal{Q}}\lambda_{\max}(\boldsymbol{R}_{qnk})$
9
10for the strongest user alignment $\boldsymbol{R}_{q^{*}nk}$ do
@@
14Compute the jamming power allocation weighting $h_{nk}=p_{q*nk}\sum_{d=1}^{N_{J}}\lambda_{q*nk,d}$
15
16Compute the jamming power scaling factors $g_{nk}=\left(P_{J}\sqrt{h_{nk}}\right)/\left(\sum_{nk}\sqrt{h_{nk}}\right)$
17
18Compute the jamming power allocation matrix $\boldsymbol{\Lambda}_{\boldsymbol{u}_{nk}}=\text{diag}\left\{\dfrac{g_{nk}\lambda_{q^{*}nk,d}}{\sum_{d=1}^{N_{J}}\lambda_{q^{*}nk,d}}\right\}_{d=1}^{N_{J}}$
19
20Compute the jamming covariance matrix $\boldsymbol{C}_{\boldsymbol{u}_{nk}}=\boldsymbol{U}_{q^{*}nk}\boldsymbol{\Lambda}_{\boldsymbol{u}_{nk}}\boldsymbol{U}_{q^{*}nk}^{H}$
@@
IV Experiments, Simulation Results and Analysis
In this section, we discuss our simulation results for applying R-SFLLM to the distributed training of LLMs. We provide insights into the sensitivity of SFL to poisoned LLM aggregations, the impact of the worst-case jammer as compared to barrage jamming, the effectiveness of sensing-assisted anti-jamming, as well as the role of different LLM architectures.
IV-A Experimental Setup
We refer to the R-SFLLM setup in Figure 2. Therein, $Q=3$ legitimate clients participate in fine-tuning BERT [38] and RoBERTa [39] base models for two distinct NLP tasks: Sequence classification (SC) and named entity recognition (NER). SC assigns a category to a sequence of words or tokens, while NER identifies named entities, such as persons or organizations within a text. For SC, the binary classification datasets SST2 [40] and QNLI [41], as well as the ternary dataset MNLI [42] are considered, while for NER the CONLL2003 [43] and WNUT17 [44] datasets are used. Each dataset is divided equally among the clients, ensuring a unique and private portion of the data. For each NLP task, the pre-trained LLMs are fine-tuned for $N_{\text{epochs}}=10$ epochs and $N_{\text{rounds}}=10$ global SFL rounds. In this setup, worst-case adversarial jamming is encountered during each uplink transmission of the LLM word embeddings in the MIMO-OFDM MAC. All participating parties employ uniform linear antenna arrays with $N_{T_{q}}=8$ and $N_{R}=16$ legitimate transmit and receive antennas. The adversary is assumed to be in the jammer-dominant regime and employs the previous worst-case strategy to maximally corrupt the LLM word embbeddings. The corresponding user and jamming DoAs are determined as
@@ IV-A Experimental Setup
The considered communication protocol employs 5G New Radio (NR) slots with $K=14$ symbols per slot and $N=64$ subcarriers. The maximum number of resource allocation blocks for each user $q$ is given by $B_{q}=\left\lfloor\frac{NK}{Q}\right\rfloor=298$.
We choose the scale parameter of the sensing-assisted R-SFLLM anti-jamming to be $\eta=10$, which is similar to previous experiments in [22] and hence much larger than the background noise $\sigma^{2}=-3\text{ dBm}$. We further resample the jamming statistics after each uplink transmission, i.e. after each training batch, to simulate movement and jamming variance. The following four scenarios are studied as benchmarks:
@@
Table I summarizes the experiment configuration and Table II shows the fine-tuning results for each scenario. We further include detailed performance plots for the aggregated global SFL model after each global round, i.e. after each $N_{\text{epochs}}$, for every dataset, base model, and scenario in Figure 3.
@@
$\phi_{H_{q,l}}$ Angle spread for legitimate users ±$\pm 10^{\circ}$
$\phi_{G,l}$ Angle spread for adversarial jammer ±$\pm 5^{\circ}$
@@
SFL Baseline Gaussian Protection No Protection
@@
Client 1 MSE $\infty$- ∞ dB -10.2 dB -7.9 dB 23.3 dB
Client 2 MSE $\infty$- ∞ dB -9.8 dB -5.9 dB 25.9 dB
Client 3 MSE $\infty$- ∞ dB -10.5 dB -3.4 dB 27.1 dB
@@
TABLE II: Fine-Tuning results for BERT and RoBERTa on SST2, QNLI, MNLI, CONLL2003, and WNUT17 datasets across different SFL training scenarios, evaluated via Accuracy (for SC) and F1 Scores (for NER).
IV-B R-SFLLM Simulation Results
IV-B1 Sequence Classification
As shown in Table II, for all three SC datasets, R-SFLLM is able to consistently safeguard the distributed training in general, leading to robust global models with classification accuracies near-identical or very close to the baselines. For instance, BERT achieves near-identical performance to the SST2 baseline with accuracies above $91\%$ for the scenarios when worst-case jamming is absent (Gaussian) and when R-SFLLM protection against it is employed (Protection). This indicates that AWGN alone does not significantly impact the word embeddings and is negligible within a statistical variance, thereby acting as a light regularizer. As for the scenario where no protection is provided (No Protection), the worst-case jammer is successful in maximally disrupting the distributed training, resulting in a global model accuracy of around $50\%$. As such, the global SFL model is no better than simple guessing, in other words, it has not been able to learn anything and subsequently has worst-case binary classification performance. This answers the question of whether worst-case jamming translates into worst-case performance positively. Moreover, each client observes near optimal performance from early epochs on, such that the global model already converges after the first global round as shown in Figure 3a. The same can be observed for MNLI and QNLI in Figures 3c and 3d. Hence, we can state that BERT is already well pre-trained on sentiment analysis, such that SC might be an easy fine-tuning task.
In comparison, RoBERTa achieves a slightly better overall baseline performance around $93\%$ for SST2, indicating a potential advantage due to its more complex architecture. However, for the scenario Protection, Figure 3b shows that the global model starts off with a slightly lower accuracy and converges to the baseline after the second global round. While negligible in this case, this may indicate that RoBERTa might be more sensitive to noisy embeddings. This is due to the fact that jamming cannot be mitigated perfectly, resulting in higher MSEs for scenario Protection as compared to Gaussian. A possible explanation is that RoBERTa does not use segment embeddings and relies solely on position and token embeddings, unlike BERT, which uses all three embedding types. Furthermore, RoBERTa’s pre-training process involves more extensive and diverse data, leading to a model that is more finely tuned to the nuances of language. This renders RoBERTa more susceptible to perturbations in the word embeddings, as it has learned to rely on subtle features that can be disrupted by noise. This noise sensitivity is more pronounced for the QNLI and partly for the MNLI dataset in Figures 3f and 3d, respectively. In particular, RoBERTa converges to the QNLI baseline only after the sixth global round.
Figure 4 further shows the QNLI performance for RoBERTa clients 1 to 3, respectively, where clients 1 and 2 approach the baseline whereas client 3 aligns more to the observed global SFL model. When further comparing the MSEs for each user in Table II, client 1 experiences the lowest MSE, while client 3 experiences the highest MSE, which is about 2.8 times higher than the one from client 1. This indicates that the developed resource allocation of the proposed protection scheme is not fair. However, this difference, while significant, does not seem to particularly affect the BERT model. This corroborates that RoBERTa is more sensitive to noisy word embeddings in general. However, since this is not necessarily observed for RoBERTa on the SST2 dataset, there likely is an additional dependence on the data distribution, such that a potentially non-independently-and-identically distributed (non-IID) data split may further decrease the performance in case of corrupted word embeddings. Thus, fine-tuning RoBERTa may result in an initially worse model if some clients underperform.
Nevertheless, R-SFLLM remains successful in mitigating the worst-case jammer, albeit after a few global rounds. This further indicates that RoBERTa makes use of the adversarial noise and robustifies over the training period, thus benefitting from the adversarial training character of R-SFLLM, in which the additional noise, if moderate, helps the model to regularize over time, thereby managing to yield a close-to-optimal global model performance after a few global rounds. This is similar to traditional adversarial training [45], however targets only word embeddings and no other layers instead.
@@ IV-B1 Sequence Classification
Figure 3: Global SFL model performance plots, evaluated after each of the $N_{\text{rounds}}$ global rounds for all four scenarios (SFL Baseline, Gaussian, No Protection, Protection) using Accuracy and F1 Scores (higher is better). Accuracy is calculated as the ratio of correctly classified sentences/words to the total number of instances, while the F1 Score is the harmonic mean of precision (the ratio of true positive observations to the total number of predicted positives) and recall (the ratio of true positive observations to the number of actual positives, i.e. the sum of true positives and false negatives).

### v3

@@ 14     Set the user scheduling to $\alpha_{qnk}=1$ for all $(n,k)\in\mathcal{I}_{q}$ and 0 otherwise
III-C Iterative Water-Filling Solution.
We have further shown in [22] that the NP-hard problem in (36) can be iteratively solved using a water-filling approach, described in Alg. 1. The proposed method adapts the original water-filling for MAC and MIMO channels [41] to incorporate user scheduling. To this end, each user $q\in\mathcal{Q}$ determines its update on the matrix $\boldsymbol{X}_{qnk}$ and computes an optimal set of the wireless system design parameters $\alpha_{qnk}$, $p_{qnk}$, $\boldsymbol{w}_{qnk}$, as outlined in steps (6)-(11), where for each user the following optimization problem is solved:
@@
Alg. 1 effectively circumvents the need to deal with the NP-hardness of the problem as the overall sum rate is indirectly maximized by maximizing the sum rate $R_{q}$ of the strongest users individually. In each iteration, we use the power iteration method for computing eigenvalues and eigenvectors, and perform water-filling for power allocation. Both of these methods are known to exhibit rapid convergence. Consequently, Alg. 1 inherits these convergence properties, which we verified in extensive experiments, where our method on average takes less than five iterations to converge. With $n_{\text{iter}}$ representing the number of iterations required for convergence, the overall complexity of the algorithm is
@@
Thus, the complexity of our anti-jamming solution increases polynomially with the number of users $Q$, hence remaining scalable for most realistic SFL deployments. We will illustrate this in our experiments in Section IV.
III-D Worst-Case Jamming Strategy.
In Proposition 2, we have seen that jamming embeddings affects the training performance. However, the impact still remains to be investigated for worst-case conditions, and in particular, how this affects the global performance in SFL after aggregating such corrupted models. To this end, we need to derive the worst-case jamming strategy, which we use to benchmark R-SFLLM. In contrast to anti-jamming, we need to find the covariance matrix $\boldsymbol{C}_{\boldsymbol{u}_{nk}}$, which minimizes $R$ instead, and pose the following adversarial objective:
@@
This is a convex semidefinite program due to the convex objective function and constraints. Thus, a global optimum can be found in polynomial time using interior-point or first-order methods [42]. However, neither one of these approaches scale efficiently for high-dimensions with a complexity of $\mathcal{O}(N_{J}^{4}N^{2}K^{2})$, or higher [43]. Thus, we require a compute-efficient alternative. To this end, we proposed a two-step approximation procedure in [22] described in Alg. 2, which consists of a prior user selection stage and a subsequent compute-efficient convex optimization. In [22], we have further shown that this worst-case jammer nullifies the sum rate, independent of the number of antennas or DoAs. Thus, we adopt this adversary as a benchmark in the subsequent experiments and investigate its impact on global model performance in SFL with LLMs to answer the question of whether worst-case jamming translates into worst-case SFL training performance.
@@
4
5for each user $q\in\mathcal{Q}$ do
@@
8
9Determine the strongest user $q^{*}=\arg\max_{q\in\mathcal{Q}}\lambda_{\max}(\boldsymbol{R}_{qnk})$
10
11for the strongest user alignment $\boldsymbol{R}_{q^{*}nk}$ do
@@
14
15Compute the jamming power allocation weighting $h_{nk}=p_{q*nk}\sum_{d=1}^{N_{J}}\lambda_{q*nk,d}$
16
17Compute the jamming power scaling factors $g_{nk}=\left(P_{J}\sqrt{h_{nk}}\right)/\left(\sum_{nk}\sqrt{h_{nk}}\right)$
18
19Compute the jamming power allocation matrix $\boldsymbol{\Lambda}_{\boldsymbol{u}_{nk}}=\text{diag}\left\{\dfrac{g_{nk}\lambda_{q^{*}nk,d}}{\sum_{d=1}^{N_{J}}\lambda_{q^{*}nk,d}}\right\}_{d=1}^{N_{J}}$
20
21Compute the jamming covariance matrix $\boldsymbol{C}_{\boldsymbol{u}_{nk}}=\boldsymbol{U}_{q^{*}nk}\boldsymbol{\Lambda}_{\boldsymbol{u}_{nk}}\boldsymbol{U}_{q^{*}nk}^{H}$
@@
IV Experiments, Simulation Results, and Analysis
In this section, we discuss our simulation results for applying R-SFLLM to SFL training with language models. We provide insights into the sensitivity of SFL to poisoned model aggregations, the impact of the worst-case jammer compared to barrage jamming, the effectiveness of our sensing-assisted anti-jamming strategy, and the influence of different model architectures. We first present detailed results for the NLP domain when using LLMs in Sec. IV-B, and then extend our analysis to VLMs in Sec. IV-C, covering two modalities across 13 diverse datasets. Finally, we present ablation results for varying the number of users $Q$ and the number of jamming antennas $N_{J}$ in R-SFLLM, demonstrating that our proposed framework is both scalable and effective.
IV-A Experimental Setup.
Figure 3: SFL setup with $Q=3$ legitimate parties, one adversarial jammer, and corresponding user and jamming DoAs, denoted by $\theta_{H_{q}}$ and $\theta_{J}$, respectively.
We refer to the R-SFLLM setup in Fig. 3, where $Q=3$ legitimate clients participate in fine-tuning LLMs and VLMs for various NLP and CV tasks and datasets.
IV-A1 NLP Experiments
We fine-tune BERT [25] and RoBERTa [26] base models for two distinct NLP tasks: Sequence classification (SC) and named entity recognition (NER). SC assigns a category to a sequence of words or tokens, while NER identifies named entities, such as persons or organizations within a text. For SC, the binary classification datasets SST2 (67k samples) [44] and QNLI (105k samples) [45], as well as the ternary dataset MNLI (393k samples) [46] are considered, while for NER CONLL2003 (14k samples) [47] and WNUT17 (3k samples) [48] are used. These datasets contain various text corpora ranging from news articles to movie reviews, thus spanning a wide range of information. For example, when fine-tuning on SST2, we perform sentiment analysis where the LLM is used to determine whether a movie review is positive or not. Each dataset is divided equally among the clients, ensuring a unique and private portion of the data. The varying dataset sizes allow us to further analyze the importance of the number of corrupted data points in SFL.
IV-A2 CV Experiments
We follow the experimental setup in [49] and fine-tune OpenAI’s CLIP ViT-B/16 VLM [27] with batch size 128 on eight different image classification datasets: MNIST (60k samples) [50], Cars (8k samples) [51], DTD (4k samples) [52], EuroSAT (21k samples) [53], GTSRB (27k samples) [54], RESISC45 (19k samples) [55], SUN397 (20k samples) [56], and SVHN (73k samples) [57]. This variety in image data distributions allows us to investigate the impact of corruptions on additional vision modalities during training.
For both NLP and CV, the pre-trained LLMs are fine-tuned for $N_{\text{epochs}}=10$ epochs and $N_{\text{rounds}}=10$ global SFL rounds.
IV-A3 Adversarial Jamming
In our SFL setup, all participating parties employ uniform linear antenna arrays with $N_{T_{q}}=8$ and $N_{R}=16$ legitimate transmit and receive antennas. The adversary is assumed to be in the jammer-dominant regime and employs the worst-case jamming strategy from Sec. III-D to maximally corrupt the embeddings during each uplink transmission in the MIMO-OFDM MAC. The corresponding user and jamming DoAs are determined as
@@ IV-A3 Adversarial Jamming
The considered communication protocol employs 5G New Radio (NR) slots with $K=14$ symbols per slot and $N=64$ subcarriers. The maximum number of resource allocation blocks for each user $q$ is given by $B_{q}=\left\lfloor\frac{NK}{Q}\right\rfloor=298$. We choose the scale parameter of the sensing-assisted R-SFLLM anti-jamming to be $\eta=10$, which is much larger than the background noise of $\sigma^{2}=-3\text{ dBm}$. This setup follows our work in [22], however, different configurations can be applied without loss of generality. We further resample the jamming statistics after each uplink transmission, i.e. after each training batch, to simulate movement and jamming variance.
The following four scenarios are studied as benchmarks:
@@
Table I summarizes the experiment configuration and Table II shows the fine-tuning results for each NLP and CV experiment scenario. For our NLP experiments, we further include detailed performance plots for the aggregated global SFL model after each global round, i.e., after each $N_{\text{epochs}}$, for every dataset, base model, and scenario in Fig. 4. For our CV experiments, we include a similar performance analysis in Fig. 10, averaged for all considered datasets. We provide relevant code and datasets as part of our GitHub repository1.
@@
$\phi_{H_{q,l}}$ Angle spread for legitimate users $\pm 10^{\circ}$
$\phi_{G,l}$ Angle spread for adversarial jammer $\pm 5^{\circ}$
@@
$b_{\text{CV\_DATASETS}}$ Batch Size for the CV experiment datasets 128
@@
IV-B R-SFLLM Simulation Results for NLP Models.
For our NLP experiments, we separately evaluate SC and NER tasks, comparing BERT and RoBERTa model architectures and examining how dataset size affects performance. In addition, we consider the scenario where only one user is jammed to better understand the impact of partial model poisoning on global SFL outcomes. Finally, we also investigate the role of the jammer by analyzing the impact of a simple barrage jammer compared to the worst-case jammer.
IV-B1 Sequence Classification
As shown in Table II, for all three SC datasets, R-SFLLM is able to consistently safeguard the distributed training in general, leading to resilient global models with classification accuracies near-identical or very close to the baselines. For instance, BERT achieves near-identical performance to the SST2 baseline with accuracies above $91\%$ for the scenarios when worst-case jamming is absent (Gaussian) and when R-SFLLM protection against it is employed (Protection). This indicates that AWGN alone does not significantly impact the embeddings and is negligible within a statistical variance, thereby acting as a light regularizer. As for the scenario where no protection is provided (No Protection), the worst-case jammer is successful in maximally disrupting the distributed training, resulting in a global model accuracy of around $50\%$. Thus, the global SFL model is no better than simple guessing, in other words, it has not been able to learn anything and subsequently has worst-case binary classification performance. This answers the question of whether worst-case jamming translates into worst-case performance positively. Moreover, each client observes near optimal performance from early epochs on, such that the global model already converges after the first global round, shown in Fig. 4a. The same can be observed for MNLI and QNLI in Fig. 4c and Fig. 4d. This demonstrates that BERT is well pre-trained such that SC is an easy fine-tuning task.
In comparison, RoBERTa achieves a slightly better overall baseline performance of around $93\%$ for SST2, indicating a potential advantage due to its more complex architecture. However, for the scenario Protection, Fig. 4b shows that the global model starts off with a slightly lower accuracy and converges to the baseline after the second global round. While negligible in this case, this indicates that RoBERTa is more sensitive to noisy embeddings. This is due to the fact that jamming cannot be mitigated perfectly, resulting in higher MSEs for scenario Protection as compared to Gaussian. In particular, RoBERTa does not use segment embeddings and relies solely on position and token embeddings, unlike BERT, which uses all three embedding types. Furthermore, RoBERTa’s pre-training process involves more extensive and diverse data, leading to a model that is more finely tuned to the nuances of language. This renders RoBERTa more susceptible to perturbations in embeddings, as it has learned to rely on subtle features that can be disrupted by noise. This noise sensitivity is more pronounced for the QNLI and partly for the larger MNLI dataset, shown in Fig. 4f and Fig. 4d, respectively, such that RoBERTa converges to the QNLI baseline only after the sixth global round.
To investigate the fairness of our resource allocation strategy, Fig. 5 shows the QNLI performance for RoBERTa clients 1 to 3, respectively, where clients 1 and 2 approach the baseline whereas client 3 aligns more to the observed global SFL model. When further comparing the MSEs for each user in Table II, client 1 experiences the lowest MSE, while client 3 experiences the highest MSE, which is about 2.8 times higher than the one from client 1. This indicates that the developed resource allocation of the proposed protection scheme is not fair. However, this difference, while significant, does not seem to particularly affect the BERT model. This corroborates that RoBERTa is more sensitive to noisy embeddings in general. However, since this is not necessarily observed for RoBERTa on the SST2 dataset, there likely is an additional dependence on the data distribution, such that a potentially non-independently-and-identically distributed (non-IID) data split may further decrease the performance in case of corrupted embeddings. Thus, fine-tuning RoBERTa may result in an initially worse model if some clients underperform.
Nevertheless, R-SFLLM remains successful in mitigating the worst-case jammer, albeit after a few global rounds. This shows that RoBERTa makes use of the adversarial noise and robustifies over the training period, thus benefitting from the adversarial training character of R-SFLLM, in which the additional noise, if moderate, helps the model to regularize over time, thereby managing to yield a close-to-optimal global model performance after a few global rounds. This is similar to traditional adversarial training [58], however targets only embeddings and no other layers instead.
Dataset / Model / MSE SFL Baseline Gaussian Protection No Protection
NLP Results
@@ IV-B1 Sequence Classification
CV Results
MNIST 98.5% 98.0% 97.5% 9.8%
Cars 86.3% 85.3% 84.1% 7.5%
DTD 81.5% 80.4% 78.6% 6.7%
EuroSAT 97.7% 97.2% 96.7% 10.8%
GTSRB 94.3% 93.8% 93.5% 8.5%
RESISC45 96.5% 96.0% 95.5% 5.3%
SUN397 78.2% 77.7% 76.9% 10.2%
SVHN 93.1% 92.6% 92.1% 7.7%
MSE
Client 1 MSE -$\infty$ dB -10.2 dB -7.9 dB 23.3 dB
Client 2 MSE -$\infty$ dB -9.8 dB -5.9 dB 25.9 dB
Client 3 MSE -$\infty$ dB -10.5 dB -3.4 dB 27.1 dB
@@
TABLE II: Fine-Tuning results for NLP and CV experiments. NLP: BERT and RoBERTa LLMs on SST2, QNLI, MNLI, CONLL2003, and WNUT17 datasets across different SFL training scenarios, evaluated via Accuracy (for SC) and F1 Scores (for NER). CV: CLIP ViT-B/16 VLM on MNIST, Cars, DTD, EuroSAT, GTSRB, RESISC45, SUN397, and SVHN image classification datasets, evaluated via Accuracy.
@@
Figure 4: Global SFL model performance plots for NLP experiments with BERT/RoBERTa, evaluated after each of the $N_{\text{rounds}}$ global rounds for all four scenarios (SFL Baseline, Gaussian, No Protection, Protection) using Accuracy and F1 Scores (higher is better). Accuracy is calculated as the ratio of correctly classified sentences/words to the total number of instances, while the F1 Score is the harmonic mean of precision (ratio of true positive observations to the total number of predicted positives) and recall (ratio of true positive observations to the number of actual positives, i.e. the sum of true positives and false negatives).